The decomposition of glucose in aqueous acid solution was found to follow first 

order. The initial concentration was found to be 0.056 M. The concentration after 

a period of 12 hours was 4.10x10-2 mol/L. 

a. Calculate reaction rate constant.

b. Calculate the quantity of glucose remaining undecomposed after 8 hours.

c. Estimate the amount of glucose lost during the period of 24 hours.


1
Expert's answer
2021-04-29T10:44:03-0400

a)


"0.041=0.056e^{-12k}\\k=0.026\ h^{-1}"

b)


"x=0.056e^{-0.026(8)}=0.0455\frac{mol}{L}"

c)


"x=0.056(1-e^{-0.026(24)})=0.026\frac{mol}{L}"